OpenStudy (anonymous):

find the sum of integers between 200 and 300 that are divisible by 7.

OpenStudy (gorv):

after 200 first number divisible by 7 is 203 next will be....210 217 224.............294 so no. are 203 210 217 224 231...........294 it is an arithmatic progression with common difference= 7....now do it @meghanamaggi

OpenStudy (anonymous):

thnku

OpenStudy (anonymous):

bt which formula to put!?

OpenStudy (gorv):

find total no of terms ...... 294=203+(n-1)*7 294-203=91=(n-1)*7 n-1=13 n=13+1=14

OpenStudy (gorv):

sum = (n/2)*(first +last term) =(14/2)*(203+294)=7*497

OpenStudy (anonymous):

thnku